Find an equation of a line containing the given pair of points. Express your answer in the form a=x, b=y, or y=mx+b
(-2,-2)(8,8)
what is an equation of the line y=

Answer by mananth(16949)

x1 y1 x2 y2
-2 -2 8 8

slope m = (y2-y1)/(x2-x1)
( 8 - -2 )/( 8 - -2 )
( 10 / 10 )
m= 1

Plug value of the slope and point ( -2,-2 ) in
Y = m x + b
-2.00 = -2 + b
b= -2 - -2
b= 0
So the equation will be
Y =x
m.ananth@hotmail.ca